# On Leakage of Code Generation Evaluation Datasets

### v1

@@ 5 Limitations
claude-2 Anthropic 0.65
claude-2.1 Anthropic 0.70
claude-3-haiku Anthropic 0.77
claude-3-sonnet Anthropic 0.74
claude-3-opus Anthropic 0.84
Command R+ Cohere 0.65
Command R Cohere 0.43
DBRX Instruct DBRX 0.60
deepseek-coder-33b-instruct DeepSeek 0.73
codellama-7b-instruct Meta 0.39
codellama-34b-instruct Meta 0.53
codellama-70b-instruct Meta 0.51
llama2-7b-chat Meta 0.17
llama2-70b-chat Meta 0.32
llama3-8b-instruct Meta 0.62
llama3-70b-instruct Meta 0.82
Mistral-7b-instruct Mistral 0.31
Mistral-7b-instruct v2 Mistral 0.42
Mistral-7b-instruct v3 Mistral 0.43
Mixtral 8x7B Mistral 0.53
mistral-small Mistral 0.63
mistral-medium Mistral 0.60
mistral-large Mistral 0.68
Mixtral 8x22B Mistral 0.73
Codestral 22B Mistral 0.82
gpt-3.5-turbo 01/25 OpenAI 0.75

### v2

@@ 5 Limitations
Mistral-7b Mistral 0.11
Mixtral 8x7B Mistral 0.17
Mixtral 8x22B Mistral 0.38
Codestral 22B Mistral 0.40
Codellama-7b-instruct Meta 0.14
Codellama-34b-instruct Meta 0.19
Llama2-7b-chat Meta 0.02
Llama2-70b-chat Meta 0.10
Llama3-8b-instruct Meta 0.27
Llama3-70b-instruct Meta 0.49
GPT-3.5-tb OpenAI 0.37
GPT-4 OpenAI 0.53
GPT-4o OpenAI 0.63
Claude-2 Anthropic 0.34
Claude-3-haiku Anthropic 0.34
Claude-3-sonnet Anthropic 0.40
Claude-3-opus Anthropic 0.54
Deepseek-coder-33b-instruct DeepSeek 0.39
DBRX Instruct DBRX 0.25
Qwen1.5-72B-Chat Qwen 0.20
Qwen1.5-110B-Chat Qwen 0.30
Qwen 2-72B Instruct Qwen 0.42
Command R+ Cohere 0.22
Command R Cohere 0.12